OpenStudy (anonymous):

Complete the left half of the graph of y=f(x) for each of the following conditions: (a) f(-x) = f(x) (b) f(-x)= -f(x) How exactly do I solve this? Isn't y=f(x) essentially y=y? I just need a little help on how to understand this problem :\

OpenStudy (anonymous):

In the case of f(-x) = f(x), the left side would just be a reflection of the right side.

OpenStudy (anonymous):

In the case of f(-x) = -f(x) you would reflect horizontally and then vertically.

OpenStudy (anonymous):

y=f(x) just means that your y coordinate will be determined by f(x).

OpenStudy (anonymous):

Ah, I see. That makes a lot more sense, thank you!